OpenStudy (anonymous):

Solve. 3x2 = –15x + 21

OpenStudy (anonymous):

*3x^2 = -15x +21

OpenStudy (turingtest):

3x^2+15x-21=0 x^2+5x-7=0 quadratic formula time with a=1, b=5, c=-7 You remember the quadratic formula, right?

OpenStudy (anonymous):

Yes i do (:

OpenStudy (turingtest):

That's all you can do here since it's not factorable. Just plug in the numbers.

OpenStudy (turingtest):

\[x=(-b \pm \sqrt{b^2-4ac})/2a\]

OpenStudy (anonymous):

Ohhh i understand now. Thank you

OpenStudy (turingtest):

no prob